# PINSAT: Parallelized Interleaving of Graph Search and Trajectory Optimization for Kinodynamic Motion Planning

### 17 Jan 2024 v1

@@ Abstract
Trajectory optimization is a widely used technique in robot motion planning for letting the dynamics and constraints on the system shape and synthesize complex behaviors. Several previous works have shown its benefits in high-dimensional continuous state spaces and under differential constraints. However, long time horizons and planning around obstacles in non-convex spaces pose challenges in guaranteeing convergence or finding optimal solutions. As a result, discrete graph search planners and sampling-based planers are preferred when facing obstacle-cluttered environments. A recently developed algorithm called INSAT effectively combines graph search in the low-dimensional subspace and trajectory optimization in the full-dimensional space for global kinodynamic planning over long horizons. Although INSAT successfully reasoned about and solved complex planning problems, the numerous expensive calls to an optimizer resulted in large planning times, thereby limiting its practical use. Inspired by the recent work on edge-based parallel graph search, we present PINSAT, which introduces systematic parallelization in INSAT to achieve lower planning times and higher success rates, while maintaining significantly lower costs over relevant baselines. We demonstrate PINSAT by evaluating it on 6 DoF kinodynamic manipulation planning with obstacles.
1 Introduction
@@
Graph search-based planning algorithms like A* and its variants (Hart, Nilsson, and Raphael 1968; Pohl 1970; Aine et al. 2016) enable robots to come up with well-reasoned long-horizon plans to achieve a given task objective (Kusnur et al. 2021; Mukherjee et al. 2021). They do so by searching over the graph that results from discretizing the state and action space. However, in robotics, several dynamically rich tasks require high-dimensional planning in the continuous space. For such domains, kinodynamic planning and trajectory optimization techniques have been developed to synthesize dynamically feasible trajectories. The existing kinodynamic algorithms achieve this by discretizing the action space to roll out trajectory primitives in the discrete or continuous state space. On the other hand, trajectory optimization techniques do not discretize the state or action space but suffer local minima, lack convergence guarantees in nonlinear settings, and struggle to reason over long horizons.
Figure 1: Filmstrip of waypoints along a kinodynamic plan generated using PINSAT for an ABB arm.
@@
An algorithm called INSAT (Natarajan, Choset, and Likhachev 2021; Natarajan et al. 2023), short for INterleaved Search And Trajectory optimization, bridges this gap for global kinodynamic planning. The idea behind INSAT was (a) to identify a low-dimensional manifold, (b) perform a search over a discrete graph embedded in this manifold, and (c) while searching the graph, utilize full-dimensional trajectory optimization to compute the cost of partial solutions found by the search. Thus every edge/action evaluation in the graph required solving at least one and potentially several instances of trajectory optimization. Though INSAT used dynamic programming to cleverly warm-start every instance of optimization with a good approximate solution, the frequent calls to the optimizer limited its practical usage.
@@
For domains where action evaluation is expensive, a parallelized planning algorithm ePA*SE (Edge-based Parallel A* for Slow Evaluations) was developed (Mukherjee, Aine, and Likhachev 2022a) that changes the basic unit of the search from state expansions to edge expansions. This decouples the evaluation of edges from the expansion of their common parent state, giving the search the flexibility to figure out what edges need to be evaluated to solve the planning problem. In this work, we employ the parallelization technique of ePA*SE to parallelize the expensive trajectory optimization step in INSAT. The resulting algorithm termed PINSAT: Parallel Interleaving of Graph Search and Trajectory Optimization, can compute dynamically feasible plans while achieving close to real-time performance. The key motivation for developing PINSAT is that if INSAT can effectively solve long-horizon, dynamically rich planning tasks, then using the ideas of ePA*SE parallelization to expedite slow edge optimizations in INSAT will result in a strictly better algorithm. We empirically demonstrate that this is indeed the case in PINSAT.
2 Related Work
Kinodynamic Motion Planning
@@ 2 Related Work
Kinodynamic planning has been a subject of extensive research, driven by the need for robots and autonomous systems to navigate complex environments while considering their dynamic constraints. This dynamic feasibility is typically achieved using one of the four different schemes below
  • Search-based kinodynamic planning involves precomputing a set of short, dynamically feasible trajectories that capture the robot’s capabilities called motion primitives. Then search-based planning algorithms like A* and its variants (Likhachev and Ferguson 2009) can be used to plan over a set of precomputed motion primitives. These search-based methods provide strong guarantees on optimality and completeness w.r.t the chosen motion primitives. However, the choice and calculation of these motion primitives that prove efficient can be challenging, particularly in high-dimensional systems.
  • Sampling-based kinodynamic methods adapt and extend classic approaches such as Probabilistic Roadmaps (PRMs) (Kavraki et al. 1996) and Rapidly-exploring Random Trees (RRTs) (LaValle et al. 1998) to handle dynamic systems (LaValle and Kuffner Jr 2001). This is achieved using dynamically feasible rollouts with random control inputs or solutions of boundary value problems within the extend operation. There are probabilistically complete and asymptotically optimal variants (Kunz and Stilman 2015; Hauser and Zhou 2016), however, empirical convergence might be tricky.
@@
  • Optimization-based planning methods can generate high-quality trajectories that are dynamically feasible and do not suffer from the curse of dimensionality in search-based methods. They formulate the motion planning problem as an optimization problem (Betts 2010) with cost functions defined for trajectory length, time, or energy consumption. After transcribing into a finite-dimensional optimization, these methods rely on the gradients of the cost function and dynamics and employ numerical optimization algorithms to find locally optimal solutions (Schulman et al. 2014; Toussaint 2017; Tassa, Mansard, and Todorov 2014). However, except for a small subset of systems (such as linear or flat systems), for most nonlinear systems these methods lack guarantees on completeness, optimality or convergence.
  • Hybrid planning methods combine two or all of the above-mentioned schemes. Search and sampling methods are combined in (Gammell, Srinivasa, and Barfoot 2015; Sakcak et al. 2019; Littlefield and Bekris 2018), sampling and optimization methods are combined in (Choudhury et al. 2016; Kamat et al. 2022; Alwala and Mukadam 2021), search, sampling and optimization methods are combined in (Ortiz-Haro et al. 2023). INSAT combined search and optimization methods and demonstrated its capability in several complex dynamical systems (Natarajan, Choset, and Likhachev 2021; Natarajan et al. 2023).
Parallel Search
@@
Several approaches parallelize sampling-based planning algorithms in which parallel processes cooperatively build a PRM (Jacobs et al. 2012) or an RRT (Devaurs, Siméon, and Cortés 2011; Ichnowski and Alterovitz 2012; Jacobs et al. 2013) by sampling states in parallel. However, in many planning domains, sampling of states is not trivial. One such class of planning domains is simulator-in-the-loop planning, which uses a physics simulator to generate successors (Liang et al. 2021). Unless the state space is simple such that the sampling distribution can be scripted, there is no principled way to sample meaningful states that can be realized in simulation.
@@
We focus on search-based planning, constructing the graph by recursively applying actions from every state. A* can be parallelized by generating successors in parallel during state expansion, but its improvement is bounded by the domain’s branching factor. Alternatively, approaches like (Irani and Shih 1986; Evett et al. 1995; Zhou and Zeng 2015) parallelize state expansions, allowing re-expansions to handle premature expansions before minimal cost. However, they may encounter an exponential number of re-expansions, especially with a weighted heuristic. In contrast, PA*SE (Phillips, Likhachev, and Koenig 2014) parallelly expands states at most once without affecting solution quality bounds. Yet, PA*SE becomes inefficient in domains with costly edge evaluations, as each thread sequentially evaluates outgoing edges. To address this, ePA*SE (Mukherjee, Aine, and Likhachev 2022a) improves PA*SE by parallelizing the search over edges. MPLP (Mukherjee, Aine, and Likhachev 2022b) achieves faster planning by lazily running the search and asynchronously evaluating edges in parallel, assuming successors can be generated without evaluating the edge. Some work focuses on parallelizing A* on GPUs (Zhou and Zeng 2015; He et al. 2021), but their SIMD execution model limits applicability to domains with simple actions sharing the same code.
@@
To the best of our knowledge, there is very little previous work on parallel kinodynamic planning. All existing approaches are sampling-based and employ a straightforward parallelization to execute the steer operation in batch—either on CPU for steering with boundary value solvers (Li et al. 2021) or on GPU for steering with neural networks in batch (Chow, Chang, and Hollinger 2023). Consequently, they inherit the same limitations as sampling-based planners, where the sampling of states can be nontrivial. Search-based methods overcome this limitation by systematically exploring the space, and INSAT & PINSAT owe their success to this characteristic.
3 Problem Formulation
@@
Let the $n$-dimensional state space of the robot be denoted by $\mathcal{X}\subseteq\mathbb{R}^{n}$. Let $\mathcal{X}^{obs}\subset\mathcal{X}$ be the obstacle space and $\mathcal{X}^{\text{free}}=\mathcal{X}\setminus\mathcal{X}^{obs}$ be the free motion planning space. For kinodynamic motion planning, we should reason about and satisfy constraints on the derivatives of position such as velocity, acceleration, etc. This boils down to including those derivatives as a part of the planning state that can quickly lead to a large intractable planning space. To cope with this let us consider a low-dimensional subspace $\mathcal{X}_{L}$ of the full-dimensional planning state space $\mathcal{X}$ such that $\mathcal{X}^{obs}\subset\mathcal{X}_{L}$. So $\mathcal{X}^{obs}\subset\mathcal{X}_{L}\subseteq\mathcal{X}$ and the auxiliary space $\mathcal{X}_{A}=\mathcal{X}\setminus\mathcal{X}_{L}$. Consider a finite graph $G=(\mathcal{V},\mathcal{E})$ defined as a set of vertices $\mathcal{V}$ and directed edges $\mathcal{E}$ embedded in this low-D space $\mathcal{X}_{L}$. Each vertex $v\in\mathcal{V}$ represents a state $\mathbf{x}\in\mathcal{X}_{L}$. An edge $e\in\mathcal{E}$ connecting two vertices $v_{1}$ and $v_{2}$ in the graph represents an action $\mathbf{a}\in\mathcal{A}$ that takes the agent from corresponding states $\mathbf{x}_{1}$ to $\mathbf{x}_{2}$. In this work, we assume that all actions are deterministic. Hence an edge $e$ can be represented as a pair $(\mathbf{x},\mathbf{a})$, where $\mathbf{x}$ is the state at which action $\mathbf{a}$ is executed. For an edge $e$, we will refer to the corresponding state and action as $e.\mathbf{x}$ and $e.\mathbf{a}$ respectively. So given a start state $\mathbf{x}^{S}$ and a goal region $\mathcal{G}$, the kinodynamic motion planning problem can be cast as the following optimization problem
@@
where $\boldsymbol{\phi}:[0,T]\rightarrow\mathcal{X}_{L}$ is a spatial trajectory that is $\gamma$-times differentiable and continuous, $w_{1}$ and $w_{2}$ trade-off the relative importance between path length and duration, and $\boldsymbol{\phi}^{(j)}(t)=\frac{d^{j}\boldsymbol{\phi}(t)}{dt^{j}}$ is the $l$-th derivative of $\boldsymbol{\phi}(t)$. The Eq. 1b is the constraint on the maximum duration of the trajectory, Eq. 1c is the boundary start and goal conditions, Eq. 1d are the system limits on the derivatives of the positional trajectory and Eq. 1e requires the trajectory to lie in $\mathcal{X}^{\text{free}}$. The minimizer of the Eq. 3 is a trajectory $\boldsymbol{\phi}_{\textbf{x}^{S}\textbf{x}^{G}}(t)$ [^1] The argument $t$ of the $\boldsymbol{\phi}(t)$ may be dropped for brevity. that connects start state $\mathbf{x}^{S}$ to the goal region $\mathcal{G}$. There is a computational budget of $N_{t}$ threads available, which can run in parallel.
4 Background
In kinodynamic planning, a great deal of focus is on the integration accuracy of the dynamic equations to obtain feasible solutions to the dynamic constraints. However, controllers in fully actuated systems, like the vast majority of commercial manipulators, can accurately track trajectories generated with smooth splines (Lau, Sprunk, and Burgard 2009) even if they are mildly inconsistent with the system’s nonlinear dynamics. Even for many other dynamical systems, parameterizing trajectories with polynomials achieves near-dynamic feasibility and proves to be highly effective. Popular techniques like direct collocation and pseudo-spectral methods parameterize trajectories with polynomials at some level. In this work, we use a piecewise polynomial widely used recently for motion planning (Liu, Lai, and Wu 2013; Kicki et al. 2023) called B-spline to represent trajectories. We will first provide some background on B-splines and highlight a few nice properties that enable us to satisfy dynamics and guarantee completeness.
4.1 B-Splines
B-splines are smooth and continuous piecewise polynomial functions made of finitely many basis polynomials called B-spline bases. A $k$-th degree B-spline basis with $m$ control points can be calculated using the Cox-de Boor recursion formula [Patrikalakis et al] as
@@ 4.1 B-Splines
where $i=0,\ldots,m$, $\frac{t-t_{i}}{t_{i+k-1}-t_{i}}$ and $\frac{t_{i+k}-t}{t_{i+k}-t_{i+1}}$ are the interpolating coefficients between $t_{i}$ and $t_{i+k}$. For $k=0$, $N_{i,0}$ is given by
$N_{i,0}(t)=\left\{\begin{array}[]{ll}1&\text{ if }t_{i}\leq t<t_{i+1}\\ 0&\text{ otherwise }\end{array}\right.$ (3)
Let us define a non-decreasing knot vector T
$\textbf{T}=\{\underbrace{t_{0},\ldots,t_{0}}_{k-\text{points}},t_{k+1},\ldots,t_{m},\underbrace{t_{f},\ldots,t_{f}}_{k-\text{points}}\}$ (4)
and the set of control points P
$\textbf{P}=\{\textbf{p}_{0},\textbf{p}_{1},\ldots,\textbf{p}_{m}\}$ (5)
where $\textbf{p}_{i}\in\mathbb{R}^{n},i=0,\ldots,m$. Then our B-spline trajectory $\boldsymbol{\phi}(t)$ can be uniquely determined by the degree of the polynomial $k$, the knot vector T, and the set of control points P called de Boor points.
@@
5 Approach
PINSAT interleaves low-D parallelized graph search with full-D trajectory optimization to combine the benefits of the former’s ability to rapidly search non-convex spaces and solve combinatorial parts of the problem and the latter’s ability to obtain a locally optimal solution not constrained to discretization. This section will outline the fundamental selection of a low-D search algorithm and the necessary modifications to make it compatible with PINSAT. Subsequently, we will elaborate on how an edge in the graph in $\mathcal{X}_{L}$ is lifted to $\mathcal{X}$ using the B-spline optimization introduced below.
5.1 B-Spline Optimization
This subsection explains the abstract optimization over the length and duration of B-splines that satisfy the obstacle, duration, derivative, and boundary constraints. In the later section, we describe how it is utilized within ePA*SE’s parallel graph search.
Decoupling Optimization over Trajectory and its Duration
It is easy to observe from Eq. 3 that there is a coupling between the trajectory $\boldsymbol{\phi}(t)$ and its duration $T$. For simultaneous optimization over the $\boldsymbol{\phi}$ and $T$, we decouple them by introducing a trajectory coordinate $u\in[0,1]$ and a mapping to convert this coordinate to the actual time, $t=\alpha(u)$ [cite]. Let the mapping $\alpha$ be monotonously increasing with $\alpha(0)=0$ and $\alpha(1)=T$ and $\boldsymbol{\phi}(t)=\boldsymbol{\phi}(\alpha(u))=\boldsymbol{\psi}(u)$. With this transformation, the derivatives of the trajectory can be derived as

### 16 Mar 2024 v2

@@ Abstract
Trajectory optimization is a widely used technique in robot motion planning for letting the dynamics of the system shape and synthesize complex behaviors. Several previous works have shown its benefits in high-dimensional continuous state spaces and under differential constraints. However, long time horizons and planning around obstacles in non-convex spaces pose challenges in guaranteeing convergence or finding optimal solutions. As a result, discrete graph search planners and sampling-based planers are preferred when facing obstacle-cluttered environments. A recently developed algorithm called INSAT effectively combines graph search in the low-dimensional subspace and trajectory optimization in the full-dimensional space for global kinodynamic planning over long horizons. Although INSAT successfully reasoned about and solved complex planning problems, the numerous expensive calls to an optimizer resulted in large planning times, thereby limiting its practical use. Inspired by the recent work on edge-based parallel graph search, we present PINSAT, which introduces systematic parallelization in INSAT to achieve lower planning times and higher success rates, while maintaining significantly lower costs over relevant baselines. We demonstrate PINSAT by evaluating it on 6 DoF kinodynamic manipulation planning with obstacles.
I Introduction
@@
Graph search-based planning algorithms like A* and its variants [1, 2, 3] enable robots to come up with well-reasoned long-horizon plans to achieve a given task objective [4, 5]. They do so by searching over the graph that results from discretizing the state and action space. However, in robotics, several dynamically rich tasks require high-dimensional planning in the continuous space. For such domains, kinodynamic planning and trajectory optimization techniques have been developed to synthesize dynamically feasible trajectories. The existing kinodynamic algorithms achieve this by discretizing the action space to roll out trajectory primitives in the discrete or continuous state space. On the other hand, trajectory optimization techniques do not discretize the state or action space but suffer local minima, lack convergence guarantees in nonlinear settings, and struggle to reason over long horizons.
Figure 1: ABB arm evading obstacles to block a ball thrown at it using kinodynamic motion produced by PINSAT.
@@
An algorithm called INSAT [6, 7], short for INterleaved Search And Trajectory optimization, bridges this gap for global kinodynamic planning. The idea behind INSAT was (a) to identify a low-dimensional manifold, (b) perform a search over a discrete graph embedded in this manifold, and (c) while searching the graph, utilize full-dimensional trajectory optimization to compute the cost of partial solutions found by the search. Thus every edge/action evaluation in the graph required solving at least one and potentially several instances of trajectory optimization. Though INSAT used dynamic programming to cleverly warm-start every instance of optimization with a good approximate solution, the frequent calls to the optimizer limited its practical usage.
@@
For domains where action evaluation is expensive, a parallelized planning algorithm ePA*SE (Edge-based Parallel A* for Slow Evaluations) was developed [8] that changes the basic unit of the search from state expansions to edge expansions. This decouples the evaluation of edges from the expansion of their common parent state, giving the search the flexibility to figure out which edges need to be evaluated to solve the planning problem. In this work, we employ the parallelization technique of ePA*SE to parallelize the expensive trajectory optimization step in INSAT. The resulting algorithm termed PINSAT: Parallel Interleaving of Graph Search and Trajectory Optimization, can compute dynamically feasible plans while achieving close to real-time performance. The key motivation for developing PINSAT is that if INSAT can effectively solve long-horizon, dynamically rich planning tasks, then using the ideas of ePA*SE parallelization to expedite slow edge optimizations in INSAT will result in a strictly better algorithm. We empirically demonstrate that this is indeed the case in PINSAT.
II Related Work
II-1 Kinodynamic Motion Planning
@@
Kinodynamic planning has been a subject of extensive research, driven by the need for robots and autonomous systems to navigate complex environments while considering their dynamic constraints. This dynamic feasibility is typically achieved using four different schemes namely (i) search-based, (ii) sampling-based, (iii) optimization-based, and (iv) hybrid methods.
@@
Search-based kinodynamic planning involves precomputing a set of short, dynamically feasible trajectories called motion primitives that capture the robot’s capabilities. Search algorithms like A* or its variants [9] are used to search over a graph wherein for any state its successor states are being computed by applying motion primitives and provide guarantees of optimality and completeness w.r.t. the chosen primitives. However, the choice and calculation of these motion primitives that are efficient can be challenging, particularly in high-dimensional systems. Sampling-based kinodynamic methods adapt and extend classic approaches such as Probabilistic Roadmaps (PRM) [10] and Rapidly Exploring Random Trees (RRT) [11] to handle dynamic systems [12]. This is achieved using dynamically feasible rollouts with random control inputs or solutions of boundary value problems within the extend operation. There are probabilistically complete and asymptotically optimal variants [13, 14], however, empirical convergence might be tricky. Optimization-based planning methods can generate high-quality trajectories that are dynamically feasible and do not suffer from the curse of dimensionality in search-based methods. They formulate the motion planning problem as an optimization problem [15] with cost functions defined for trajectory length, time, or energy consumption. After transcribing into a finite-dimensional optimization, these methods rely on the gradients of the cost function and dynamics and employ numerical optimization algorithms to find locally optimal solutions [16, 17, 18]. However, except for a small subset of systems (such as linear or flat systems), for most nonlinear systems these methods lack guarantees on completeness, optimality or convergence. Hybrid planning methods combine two or all of the schemes mentioned above. Search and sampling methods are combined in [19, 20, 21], sampling and optimization methods are combined in [22, 23, 24], search, sampling, and optimization methods are combined in [25]. INSAT combined search and optimization methods and demonstrated its capability in several complex dynamical systems [6, 7].
II-2 Parallel Search
@@
Several approaches parallelize sampling-based planning algorithms in which parallel processes cooperatively build a PRM [26] or an RRT [27, 28, 29] by sampling states in parallel. However, in many planning domains, sampling of states is not trivial. One such class of planning domains is simulator-in-the-loop planning, which uses a physics simulator to generate successors [30]. Unless the state space is simple, such that the sampling distribution can be scripted, there is no principled way to sample meaningful states that can be realized in simulation.
@@
Search-based methods like A* can be parallelized by generating successors concurrently during state expansion but are limited by the domain’s branching factor. Alternatively, approaches like [31, 32, 33] parallelize state expansions, allowing re-expansions to handle premature expansions before minimal cost, but may face exponential re-expansions, especially with weighted heuristics. On the contrary, PA*SE [34] parallelly expands states at most once without affecting solution quality bounds. However, PA*SE is inefficient in domains with costly edge evaluations as each thread sequentially evaluates the outgoing edges. To address this, ePA*SE [8] improves PA*SE by parallelizing edge search. MPLP [35] achieves faster planning by lazily running the search and asynchronously evaluating edges in parallel, assuming that successors can be generated without evaluating the edge. Some work focuses on parallelizing A* on GPUs [33, 36], but their SIMD execution model limits applicability to domains with simple actions sharing the same code.
@@
To the best of our knowledge, there is very little previous work on parallel kinodynamic planning. All existing approaches are sampling-based and employ a straightforward parallelization to execute the steer operation in batch—either on CPU for steering with boundary value solvers [37] or on the GPU for steering with neural networks in batch [38]. Consequently, they inherit the same limitations as sampling-based planners, where the sampling of states can be nontrivial. Search-based methods overcome this limitation by systematically exploring the space, and INSAT & PINSAT owe their success to this characteristic.
III Problem Formulation
@@
Let the $n$-dimensional state space of the robot be denoted by $\mathcal{X}\subseteq\mathbb{R}^{n}$. Let $\mathcal{X}^{obs}\subset\mathcal{X}$ be the obstacle space and $\mathcal{X}^{\text{free}}=\mathcal{X}\setminus\mathcal{X}^{obs}$ be the free motion planning space. For kinodynamic motion planning, we should reason about and satisfy constraints on the derivatives of position such as velocity, acceleration, etc. This boils down to including those derivatives as a part of the planning state and can quickly lead to a large intractable planning space. To cope with this let us consider a low-dimensional (low-D) space $\mathcal{X}_{L}$ and an auxiliary space $\mathcal{X}_{A}$ whose Cartesian product forms the full-dimensional (full-D) space $\mathcal{X}=\mathcal{X}_{A}\times\mathcal{X}_{L}$. Let the low-D space be such that $\mathcal{X}^{obs}\subset\mathcal{X}_{L}$. Consider a finite graph $G=(\mathcal{V},\mathcal{E})$ defined as a set of vertices $\mathcal{V}$ and directed edges $\mathcal{E}$ embedded in this low-D space $\mathcal{X}_{L}$. Each vertex $v\in\mathcal{V}$ represents a state $\mathbf{x}\in\mathcal{X}_{L}$. An edge $e\in\mathcal{E}$ connecting two vertices $v_{1}$ and $v_{2}$ in the graph represents an action $\mathbf{a}\in\mathcal{A}$ that takes the agent from corresponding states $\mathbf{x}_{1}$ to $\mathbf{x}_{2}$. In this work, we assume that all actions are deterministic. Hence an edge $e$ can be represented as a pair $(\mathbf{x},\mathbf{a})$, where $\mathbf{x}$ is the state at which action $\mathbf{a}$ is executed. For an edge $e$, we will refer to the corresponding state and action as $e.\mathbf{x}$ and $e.\mathbf{a}$ respectively. So given a start state $\mathbf{x}^{S}$ and a goal region $\mathcal{G}$, the kinodynamic motion planning problem can be cast as the following optimization problem
@@
where $\boldsymbol{\phi}:[0,T]\rightarrow\mathcal{X}_{L}$ is a spatial trajectory that is $\gamma$-times differentiable and therefore continuous, $w_{1}$ and $w_{2}$ trade-off the relative importance between path length and duration, and $\boldsymbol{\phi}^{(j)}(t)=\frac{d^{j}\boldsymbol{\phi}(t)}{dt^{j}}$ is the $j$-th derivative of $\boldsymbol{\phi}(t)$. The Eq. 1b is the constraint on the maximum duration of the trajectory, Eq. 1c is the boundary start and goal conditions, Eq. 1d is the element-wise system limit on the derivatives of the positional trajectory, Eq. 1e requires the trajectory to lie in $\mathcal{X}^{\text{free}}$ and Eq. 1f imposes the degree of continuity on the trajectory. The minimizer of the Eq. III is a trajectory $\boldsymbol{\phi}_{\textbf{x}^{S}\textbf{x}^{G}}(t)$ [^1] The argument $t$ of the $\boldsymbol{\phi}(t)$ may be dropped for brevity. that connects start state $\mathbf{x}^{S}$ to the goal region $\mathcal{G}$. There is a computational budget of $N_{t}$ threads available, which can run in parallel.
IV Background
In kinodynamic planning, a great deal of focus is on the integration accuracy of the equations of motion to obtain feasible solutions to the dynamic constraints. However, fully controllable systems, like the vast majority of commercial manipulators, can accurately track trajectories generated with smooth splines [39] even if they are mildly inconsistent with the system’s nonlinear dynamics. Even for many other dynamical systems, parameterizing trajectories with polynomials achieves near-dynamic feasibility and proves to be highly effective. Popular techniques like direct collocation and pseudo-spectral methods parameterize trajectories with polynomials at some level. In this work, we use a piecewise polynomial widely used recently for motion planning [40, 41] called B-spline to represent trajectories. We will first provide some background on B-splines and highlight a few nice properties that enable us to satisfy dynamics and guarantee completeness.
IV-A B-Splines
B-splines are smooth and continuous piecewise polynomial functions made of finitely many basis polynomials called B-spline bases. A $k$-th degree B-spline basis with $m$ control points can be calculated using the Cox-de Boor recursion formula [42] as
@@ IV-A B-Splines
where $i=0,\ldots,m$, $\frac{t-t_{i}}{t_{i+k-1}-t_{i}}$ and $\frac{t_{i+k}-t}{t_{i+k}-t_{i+1}}$ are the interpolating coefficients between $t_{i}$ and $t_{i+k}$. For $k=0$, $N_{i,0}=1$ if $t_{i}\leq t<t_{i+1}$ and 0 otherwise.
Let us define a non-decreasing knot vector T and the set of control points P called de Boor points
$\textbf{T}=\{\underbrace{t_{0},\ldots,t_{0}}_{k-\text{points}},t_{k+1},\ldots,t_{m},\underbrace{t_{f},\ldots,t_{f}}_{k-\text{points}}\}$ (3a)
$\textbf{P}=\{\textbf{p}_{0},\textbf{p}_{1},\ldots,\textbf{p}_{m}\}$ (3b)
where $\textbf{p}_{i}\in\mathbb{R}^{n},i=0,\ldots,m$. Then our B-spline trajectory $\boldsymbol{\phi}(t)$ can be uniquely determined by T, P and the degree of the polynomial $k$
@@
V Approach
PINSAT interleaves parallelized discrete graph search in low-D with trajectory optimization in full-D to combine the benefits of the former’s ability to search non-convex spaces and solve combinatorial parts of the problem and the latter’s ability to obtain a locally optimal solution not constrained to discretization. In the following subsection, we will justify the choice of the low-D search algorithm. Subsequently, we will explain how an edge in the graph in $\mathcal{X}_{L}$ is lifted to $\mathcal{X}$ using B-spline optimization and the necessary modifications to the low-D search to make it compatible with PINSAT.
V-A Low Dimensional Graph Search
The low-D search runs w-eA* [8] with variations that allow it to be interleaved with full-D trajectory optimization. Using w-eA* instead of wA* (like in INSAT) for low-D search provides a systematic framework to parallelize the optimization of expensive trajectory with its parallelized variant w-ePA*SE. This is because, unlike in wA*, where the basic operation of the search loop is the expansion of a state, in w-eA*, the basic operation is expansion of an edge.
V-A1 w-eA*
V-B B-Spline Optimization
This subsection explains the abstract optimization over the length and duration of B-splines that satisfy the obstacle, duration, derivative, and boundary constraints. In the later section, we describe how it is utilized within PINSAT’s parallel low-D graph search.
V-B1 Decoupling Optimization over Trajectory and its Duration
It is easy to observe from Eq. III that there is a coupling between the trajectory $\boldsymbol{\phi}(t)$ and its duration $T$. For simultaneous optimization over the $\boldsymbol{\phi}$ and $T$, we decouple them by introducing a trajectory coordinate $u\in[0,1]$ and a mapping to convert this coordinate to the actual time, $t=\alpha(u)$ [39]. Let the mapping $\alpha$ be monotonically increasing with $\alpha(0)=0$ and $\alpha(1)=T$ and $\boldsymbol{\phi}(t)=\boldsymbol{\phi}(\alpha(u))=(\boldsymbol{\phi}\circ\alpha)(u)=\boldsymbol{\psi}(u)$. The derivatives of the trajectory are